# A solution to Banach conjecture

### v5

@@ 3. Proof of the main theorem
First, we introduce a function $\Gamma:G(n,n-2)\to\mathbb{R}_{\geq 0}$ defined by
$\Gamma(H):=\sup_{\theta\in\mathbb{S}^{n-1}\cap H\atop\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}}\Big\{\Big|\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}-\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}\Big|\Big\}.$
It is clear that Condition (1) is equivalent to $\Gamma(H)=0$ for all $H\in G(n,n-2)$. And hence, We once more partition G(n,n2)G(n,n-2) into two subsets, $B_{1}$ and $B_{2}$. Define
$B_{1}:=\{H\in G(n,n-2):\Gamma(H)=0\},\quad B_{2}:=\{H\in G(n,n-2):\Gamma(H)>0\}.$
Our goal is to show that if $B_{2}\neq\emptyset$, $B_{2}=G(n,n-2)$.
Moreover, $\Gamma$ is continuous because $\Phi(\xi)$ is continuous and $J(K\cap\xi_{0}^{\perp})$ is invariant under $G_{K\cap\xi_{0}^{\perp}}$. Furthermore, $\Gamma(H_{1})=\Gamma(H_{2})$ whenever there exists a linear transformation between $K\cap H_{1}$ and $K\cap H_{2}$, as implied by Lemma 3.1. Now we need to deal with the linear transformation inside G(n,n2)G(n,n-2).
Observe that, for any $H\in G(n,n-2)$ and any $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$, there exist two linear maps $\phi^{-1}_{\xi},\phi^{-1}_{\eta}$ that send $H$ back to $\xi_{0}^{\perp}$. Depending on the relative positions of $\phi^{-1}_{\xi}(H)$ and $\phi^{-1}_{\eta}(H)$, we divide G(n,n2)G(n,n-2) into two subsets, denoted $A_{1}$ and $A_{2}$.
The set $A_{1}$ consists of all $H$ such that, for every $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$,
$\phi^{-1}_{\eta}(K\cap H)=g\circ\phi^{-1}_{\xi}(K\cap H)$
for some $g\in G_{K\cap\xi_{0}^{\perp}}$.
The set $A_{2}$ consists of all $H\in G(n-2)$ for which there exists a pair of vectors $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$ such that
$\phi^{-1}_{\eta}(K\cap H)\neq g\circ\phi^{-1}_{\xi}(K\cap H)$
for every $g\in G_{K\cap\xi_{0}^{\perp}}$. It is clear that $A_{1}\cap A_{2}=\emptyset$.
Note that, for $H\in A_{1}$ and $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$, we have
$\phi^{-1}_{\eta}(K\cap H)=g\circ\phi^{-1}_{\xi}(K\cap H)$
for some $g\in G_{K\cap\xi_{0}^{\perp}}$. Then
$\phi^{-1}_{\eta}(K\cap H)\cap J(K\cap\xi_{0}^{\perp})=g\circ\phi^{-1}_{\xi}(K\cap H)\cap J(K\cap\xi_{0}^{\perp}),$
which implies
$\phi^{-1}_{\eta}(K\cap H)\cap J(K\cap\xi_{0}^{\perp})=g(\phi^{-1}_{\xi}(K\cap H)\cap g^{-1}(J(K\cap\xi_{0}^{\perp}))),$
where $g^{-1}(J(K\cap\xi_{0}^{\perp}))=J(K\cap\xi_{0}^{\perp})$, and thus, we have
$\phi^{-1}_{\eta}(K\cap H)\cap J(K\cap\xi_{0}^{\perp})=g(\phi^{-1}_{\xi}(K\cap H)\cap J(K\cap\xi_{0}^{\perp})),$
and hence,
$\phi^{-1}_{\eta}(K\cap H\cap\phi_{\eta}(J(K\cap\xi_{0}^{\perp})))=g(\phi^{-1}_{\xi}(K\cap H\cap\phi_{\xi}(J(K\cap\xi_{0}^{\perp})))),$
that is,
$\phi^{-1}_{\eta}(\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))\cap H)=g\circ\phi^{-1}_{\xi}(\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))\cap H).$
Therefore, as a result of $\phi^{-1}_{\eta}(K\cap H)=g\circ\phi^{-1}_{\xi}(K\cap H)$, we obtain
$\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))\cap H=\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))\cap H$
which means
$\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta),$
for any $\theta\in\mathbb{S}^{n-1}\cap H$. And thus, $\Gamma(H)=0$, so $A_{1}\subseteq B_{1}$, which implies $B_{2}\subseteq A_{2}$.
We now focus on the set $B_{2}$. For any $H\in B_{2}$, and therefore $H\in A_{2}$, and for any $\xi,\eta\in\mathbb{S}^{n-1}\cap H^{\perp}$ with $\xi\neq\eta$, there exists a linear transformation
$\phi_{\eta}\circ g\circ\phi^{-1}_{\xi}(K\cap H)=\phi_{\eta}\circ g\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})=K\cap\eta^{\perp}\cap\phi_{\eta}\circ g\circ\phi^{-1}_{\xi}(\eta^{\perp}).$
Moreover, since $H\in A_{2}$, there is some $\eta,\xi\in\mathbb{S}^{n-1}\cap H^{\perp}$ such that
$\phi^{-1}_{\eta}(K\cap\xi^{\perp}\cap\eta^{\perp})\neq g\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})$
for every $g\in G_{K\cap\xi_{0}^{\perp}}$. In particular, we have $\phi_{\eta}\circ g\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})\neq K\cap\xi^{\perp}\cap\eta^{\perp}$, and by the continuity of $\Phi$, there is a path joining $\phi_{\eta}\circ g_{\eta}\circ\phi^{-1}_{\xi}(K\cap\xi^{\perp}\cap\eta^{\perp})$ to $K\cap\xi^{\perp}\cap\eta^{\perp}$ by $K\cap\zeta^{\perp}\cap\phi_{\zeta}\circ g_{\zeta}\circ\phi^{-1}_{\xi}(\zeta^{\perp})$ with $\zeta$ lying on the arc between $\xi$ and $\eta$.
Thus, as $\zeta$ moves past $\xi$, there are two possibilities: either $K\cap\eta^{\perp}\cap\phi_{\zeta}\circ g_{\zeta}\circ\phi^{-1}_{\xi}(\zeta^{\perp})$ returns to $K\cap\xi^{\perp}\cap\zeta^{\perp}$ exactly $m$ times, giving rise to an $m$-circle, or $K\cap\eta^{\perp}\cap\phi_{\zeta}\circ g_{\zeta}\circ\phi^{-1}_{\xi}(\eta^{\perp})$ arrives at $K\cap\xi^{\perp}\cap\zeta^{\perp}$, meaning that $\lim_{\zeta\to-\xi}\phi_{\zeta}\circ g_{\zeta}\circ\phi^{-1}_{\xi}(\theta)=-\theta$ for $\theta\in\xi^{\perp}$.
Owing to the central symmetry of $K$, we introduce the family $\mathcal{L}(H)$ defined by
$\mathcal{L}(H):=\{\zeta^{\perp}\cap\phi_{\zeta}\circ g_{\zeta}\circ\phi^{-1}_{\xi}(\zeta^{\perp}):\zeta,\xi\in\mathbb{S}^{n-1}\cap H^{\perp}\}.$
Consequently, $\mathcal{L}(H)\cap\mathbb{S}^{n-1}$ contains either a $2m$-loop or a single large loop passing through the antipodal point. Moreover, for any $H_{1},H_{2}\in\mathcal{L}(H)\cap G(n,n-2)$, the sections $K\cap H_{1}$ and $K\cap H_{2}$ are linearly dependent.
We now assume that $B_{2}$ is nonempty, which implies that the maximum value of $\Gamma(H)$, denoted by $M$, is nonzero. Hence, by the preceding argument and the fact that $\Gamma$ is invariant under linear transformations, the level set
$\{H\in G(n,n-2):\Gamma(H)=M\}$
definitely contains $\mathcal{L}(H)$ for some $H$ satisfying $\Gamma(H)=M$, where $\mathcal{L}(H)$ either contains an (n1)(n-1)-dimensional manifold with antipodal points, or it contains a subset of the form $H\times\infty$, where $\infty$ is the shape of a figure eight.
Observe that, by the compactness of G(n,n2)G(n,n-2), there exists a sufficiently small $\epsilon>0$ such that for every $c\in(M-\epsilon,M)$, the set
$\{H\in G(n,n-2):\Gamma(H)=c\}$
cannot contain $H\times\infty$, and therefore it must contain an (n1)(n-1)-dimensional manifold with antipodal points. Consequently, by the continuity of $\Gamma$, the level set
$\{H\in G(n,n-2):\Gamma(H)=M\}$
also contains an (n1)(n-1)-dimensional manifold with antipodal points. This implies that
$\{H\in G(n,n-2):\Gamma(H)=c\}\mbox{ and }\{H\in G(n,n-2):\Gamma(H)=M\}$
are the same set, and therefore
$\{H\in G(n,n-2):\Gamma(H)=M\}=G(n,n-2).$
We now show that if $B_{2}\neq\emptyset$, then in fact $B_{2}=G(n,n-2)$, which implies $A_{2}=G(n,n-2)$. In other words, for any $H\in G(n,n-2)$, the set $\mathcal{H}$ either contains an (n1)(n-1)-dimensional manifold with antipodal points, or it contains a subset of the form $H\times\infty$, such that for all $H_{1},H_{2}\in\mathcal{L}(H)\cap G(n,n-2)$, the sections $K\cap H_{1}$ and $K\cap H_{2}$ are linearly dependent.
We now introduce another family $\mathcal{F}(H)$ defined by
$\mathcal{F}(H):=\{\tilde{H}\in G(n,n-2):K\cap\tilde{H}\ \text{and}\ K\cap H\ \text{are linearly dependent}\}.$
Observe that $\mathcal{L}(H)\subseteq\mathcal{F}(H)$. We then assert that there can be only finitely many pairwise disjoint sets $\{H_{i}\in G(n,n-2)\}_{i=1}^{m}$, each containing a subset of the form $H\times\infty$, such that
$\mathcal{F}(H_{i})\cap\mathcal{F}(H_{j})=\emptyset$
for $i\neq j$. Suppose, to the contrary, that there are infinitely many pairwise disjoint sets $\{H_{i}\}_{i=1}^{\infty}$. Then, for each $i$, the set $H_{i+1}\times\infty$ is contained in $C_{i}$, one of the connected components of the complement determined by $H_{i}\times\infty$. By the compactness of G(n,n2)G(n,n-2), both sequences $\{C_{i}\}$ and $\{H_{i}\}$ converge to some limit $\bar{H}$. However, for all sufficiently large $i\geq I$, we have
@@
which implies
$\mathcal{L}(\bar{H})\cap\mathcal{L}(H_{i})\neq\emptyset$
for all $i\geq I$. This yields a contradiction, since it shows that there must exist a linear transformation relating the $H_{i}$ for all $i\geq I$.
We now choose $H_{0}$ such that $\mathcal{L}(H_{0})$ contains an (n1)(n-1)-dimensional manifold with antipodal points. For any other $H$ for which $\mathcal{L}(H)$ also contains an (n1)(n-1)-dimensional manifold with antipodal points, we obtain
$\mathcal{L}(H_{0})\cap\mathcal{L}(H)\neq\emptyset,$
and consequently,
$\mathcal{L}(H)\subseteq\mathcal{F}(H_{0}).$
If for some $i\in[1,m]$ we have
$\mathcal{F}(H_{0})\cap\mathcal{F}(H_{i})\neq\emptyset,$
then it follows that
$\mathcal{F}(H_{0})=\mathcal{F}(H_{i}).$
Otherwise, for every $i\in[1,m]$,
$\mathcal{F}(H_{0})\cap\mathcal{F}(H_{i})=\emptyset.$
Consequently, G(n,n2)G(n,n-2) is partitioned into $m$ or $m+1$ pairwise disjoint closed sets, and thus, by the connectedness of G(n,n2)G(n,n-2), we conclude that all (n2)(n-2)-dimensional sections of $K$ are linearly dependent. Consequently, by Gromov’s theorem [7], $K$ must be an ellipsoid. ∎
In particular, Condition (1) in Lemma 3.3 gives
$\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{\phi_{\eta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=\rho_{E}(\theta).$
And thus,
$E\cap\xi^{\perp}=\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))=J(K\cap\xi^{\perp}).$
Now we can show the following theorem.
Theorem 3.2.
Let $K\subseteq\mathbb{R}^{n}$ for $n\geq 3$ be an origin-symmetric convex body. Fixed a $\xi_{0}\in\mathbb{S}^{n-1}$; if for any $\xi\in\mathbb{S}^{n-1}$ there exists a linear transformation $\phi_{\xi}\in GL(n)$ such that $K\cap\xi^{\perp}=\phi_{\xi}(K\cap\xi_{0}^{\perp})$, then the John ellipsoid $J(K\cap\xi_{0}^{\perp})$ can form an origin-symmetric star body $E$ by
$E:=\bigcup_{\xi\in\mathbb{S}^{n-1}}\phi_{\xi}(J(K\cap\xi_{0}^{\perp})).$
Moreover, $E\cap\xi^{\perp}=J(K\cap\xi^{\perp})$ for any $\xi\in\mathbb{S}^{n-1}$ and $E=J(K)$.
Proof.
By Lemmas 3.2 and 3.3, either $K$ itself is an ellipsoid with $E=K$, or $E$ is an origin-symmetric star body satisfying $E\cap\xi^{\perp}=J(K\cap\xi^{\perp})$. In the latter case, Theorem 3.1 implies that $E$ is an ellipsoid.
Now take an arbitrary ellipsoid $\tilde{E}$ contained in $K$. For every $\xi\in\mathbb{S}^{n-1}$, we have the inequality $|\tilde{E}\cap\xi^{\perp}|\leq|E\cap\xi^{\perp}|$, because $E\cap\xi^{\perp}$ is the John ellipsoid of the section $K\cap\xi^{\perp}$. Since the ellipsoid is an intersection body, the theorem of Koldobsky and Zhang (see [9] and [16]) implies that $|\tilde{E}|\leq|E|$. Consequently, $E$ is indeed the John ellipsoid of $K$. ∎
Proof of the Theorem 1.1.
By Theorem 3.2, the body $K$ contains a John ellipsoid $E$ such that, for every $\xi\in\mathbb{S}^{n-1}$, the section $E\cap\xi^{\perp}$ is the John ellipsoid of the section $K\cap\xi^{\perp}$. By applying an affine transformation, we can move $K$ into John’s position, denoted by $\tilde{K}$, in which $E$ is transformed into a Euclidean ball $B$. Since the John ellipsoid is invariant under linear transformation, the set $\tilde{K}\cap\xi^{\perp}$ has $B\cap\xi^{\perp}$ as its John ellipsoid. Consequently, $\phi_{\xi}$ must be a congruence transformation; otherwise, the John ellipsoid of $\phi_{\xi}(\tilde{K}\cap\xi_{0}^{\perp})$ would be $\phi_{\xi}(B\cap\xi^{\perp})$, which differs from $B\cap\xi^{\perp}$.
Hence, $\tilde{K}\cap\xi^{\perp}$ has the same volume for every $\xi\in\mathbb{S}^{n-1}$, and therefore, by the Minkowski–Funk theorem, $\tilde{K}$ must be a ball, which in turn implies that $K$ is an ellipsoid. ∎
4. Acknowledgement
The crucial idea emerged when the author visited the Tianyuan Mathematical Research Center in Kunming.

### v8

@@ 3. Proof of the main theorem
By Proposition 3.1 (3), there exists some $t_{0}\in[0,1]$ such that $\Lambda_{t_{0}}=\mathbb{S}^{n-1}$. Hence, for every $\theta\in\mathbb{S}^{n-1}$, there is a $\zeta\in\mathbb{S}^{n-1}\cap\theta^{\perp}$ satisfying
$\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)\|\theta\|_{K}=t_{0}.$
Consequently,
$\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=t_{0}\rho_{K}(\theta).$
Using the definitions of $E_{t_{0}}$ and $\tilde{E}_{t_{0}}$, we obtain
$\rho_{E_{t_{0}}}(\theta)$ $=\sup\big\{\min\{\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta),t_{0}\rho_{K}(\theta)\}:\ \xi\in\mathbb{S}^{n-1}\cap\theta^{\perp}\big\}$
$=\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=t_{0}\rho_{K}(\theta).$
Similarly,
$\rho_{\tilde{E}_{t_{0}}}(\theta)$ $=\inf\big\{\max\{\rho_{\phi_{\xi}(J(K\cap\xi_{0}^{\perp}))}(\theta),t_{0}\rho_{K}(\theta)\}:\ \xi\in\mathbb{S}^{n-1}\cap\theta^{\perp}\big\}$
$=\rho_{\phi_{\zeta}(J(K\cap\xi_{0}^{\perp}))}(\theta)=t_{0}\rho_{K}(\theta).$
Thus, we conclude that
$t_{0}K=E_{t_{0}}=\tilde{E}_{t_{0}}.$
Now we can prove the main theorem.
Proof of the Theorem 1.1.
By Lemma 3.3, there exists $t_{0}\in[0,1]$ such that
$t_{0}K=E_{t_{0}}=\tilde{E}_{t_{0}}.$
On the one hand, we observe that
$E_{t}$ $=t_{0}K$
$=\bigcup_{\xi\in\mathbb{S}^{n-1}}t_{0}\phi_{\xi}\big(K\cap\xi_{0}^{\perp}\big)$
$=\bigcup_{\xi\in\mathbb{S}^{n-1}}\phi_{\xi}\big(t_{0}(K\cap\xi_{0}^{\perp})\big)$
@@
[…]
from which it follows that
[…]
On the other hand, we obtain
$t_{0}K$ $=\tilde{E}_{t}$
$=\bigcap_{\xi\in\mathbb{S}^{n-1}}\phi_{\xi}\big(J(K\cap\xi_{0}^{\perp})\cup t_{0}(K\cap\xi_{0}^{\perp})\big)$
[…]
which implies that
$J(K\cap\xi_{0}^{\perp})\backslash t_{0}(K\cap\xi_{0}^{\perp})=\emptyset.$
Consequently,
$J(K\cap\xi_{0}^{\perp})=t_{0}(K\cap\xi_{0}^{\perp}),$
so $K\cap\xi_{0}^{\perp}$ is an ellipsoid, and by Theorem 3.1 this implies that $K$ itself is an ellipsoid. ∎